OpenStudy (anonymous):

-19x^3-12x^2+16x=0 find the real solutions of the equation by graphing

OpenStudy (p0sitr0n):

like doing the graph of the function by derivating and finding critical points?

OpenStudy (anonymous):

how do you figure out or find the solution such as -1.29 -1.29,0.65 0,1.29,-0.65 0,-1.29,0.65

OpenStudy (p0sitr0n):

-19x^3-12x^2+16x=0 -19x^2-12x+16=0 use quadratic formula or factor out to find roots

OpenStudy (p0sitr0n):

a=-19 b=-12 c=16 \[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a }\]

OpenStudy (anonymous):

x=-12+-(-12)^2-4(-19)(16)/2(-19)

OpenStudy (p0sitr0n):

actually, it will begin by x=12+-, because -b= -(-12)=12

OpenStudy (p0sitr0n):

and theres a square root

OpenStudy (anonymous):

12+-144-76(64)/-38

OpenStudy (p0sitr0n):

\[x=\frac{ 12\pm \sqrt{144+4\times16\times19} }{ -2\times19}\]

OpenStudy (p0sitr0n):

x=-1.29, x=0 and x=0.65

OpenStudy (p0sitr0n):

dont forget the x=0, because at the beginning, you factored by x and then assumed either x = 0 or the other multiple is 0

OpenStudy (anonymous):

but how did you get -1.29 and 0.65

OpenStudy (p0sitr0n):

i just calculated the number under the square root. then, it says +-, so one time you solve the equation with a +, and then another time with a -. gives 2 answers, plus the x=0

OpenStudy (anonymous):

okay thats what i didnt do thanks so much for the help

OpenStudy (p0sitr0n):

youre welcome